# Near-Optimal Channel Estimation for Dense Array Systems

### v1

@@ Appendix I Proof of Corollary 1
where ${\hat{\bf P}}={\rm diag}\left(\sqrt{{\hat{p}}_{1}},\cdots,\sqrt{{\hat{p}}_{M}}\right)$ is the power allocated to the $M$ eigenvectors of $\hat{\bf\Sigma}_{\bf h}$ and ${\hat{\bf N}}={\rm diag}\left({{\hat{n}}_{1}},\cdots,{{\hat{n}}_{M}}\right)$ is the pilot-reuse-frequency matrix, as defined in (91). It is clear from (K) that the terms ${\bf W}{\bf W}^{H}$ in the three cases share a general form ${\bf W}{\bf W}^{H}={\bf{U}}{\bm{\Psi}}{\bf{U}}^{H}$ wherein ${\bm{\Psi}}={\rm diag}\left(\psi_{1},\cdots,\psi_{M}\right)$ is a diagonal matrix. Therefore, the proof of Lemma 8 can be obtained by replacing ${\bm{\Psi}}$ with the corresponding terms ${\bf{\hat{P}}}{{{\bf{\hat{P}}}}^{H}}$, ${\bf{\hat{N}}}$, and ${{Q\over M}{{\bf{I}}_{M}}}$,respectively. It allows us focus on the general proof based on the common expression ${\bf{U}}{\bm{\Psi}}{\bf{U}}^{H}$, which is given as follows.
@@
where the second equality holds since $\lambda_{m}=0$ for all $m>K$. According to (K), for the water-filling estimator, ice-filling estimator, and random estimator, we can replace the $\psi_{m}$ in (K) with ${\hat{p}}_{m}$, ${\hat{n}}_{m}$, and $\frac{Q}{M}$ to obtain ${{\hat{\delta}}_{{\rm{wf}}}}$, ${{\hat{\delta}}_{{\rm{if}}}}$, and ${{\hat{\delta}}_{{\rm{rnd}}}}$, respectively. According to (K), it is obvious that $\hat{\delta}\mathop{=}\limits^{Q\to+\infty}{\cal O}\left({{\sigma^{2}}{M^{2}}{Q^{-1}}}\right)$ holds for the three estimators. This completes the proof.

### v2

@@ Appendix I Proof of Lemma 7
where ${\hat{\bf P}}={\rm diag}\left(\sqrt{{\hat{p}}_{1}},\cdots,\sqrt{{\hat{p}}_{M}}\right)$ is the power allocated to the $M$ eigenvectors of $\hat{\bf\Sigma}_{\bf h}$ and ${\hat{\bf N}}={\rm diag}\left({{\hat{n}}_{1}},\cdots,{{\hat{n}}_{M}}\right)$ is the pilot-reuse-frequency matrix, as defined in (75). It is clear from (88) that the terms ${\bf W}{\bf W}^{H}$ in the three cases share a general form ${\bf W}{\bf W}^{H}={\bf{U}}{\bm{\Psi}}{\bf{U}}^{H}$ wherein ${\bm{\Psi}}={\rm diag}\left(\psi_{1},\cdots,\psi_{M}\right)$ is a diagonal matrix. Therefore, the proof of Lemma 7 can be obtained by replacing ${\bm{\Psi}}$ with the corresponding terms ${\bf{\hat{P}}}{{{\bf{\hat{P}}}}^{H}}$, ${\bf{\hat{N}}}$, and ${{Q\over M}{{\bf{I}}_{M}}}$,respectively. It allows us focus on the general proof based on the common expression ${\bf{U}}{\bm{\Psi}}{\bf{U}}^{H}$, which is given as follows.
@@
where the second equality holds since $\lambda_{m}=0$ for all $m>K$. According to (88), for the water-filling estimator, ice-filling estimator, and random estimator, we can replace the $\psi_{m}$ in (I) with ${\hat{p}}_{m}$, ${\hat{n}}_{m}$, and $\frac{Q}{M}$ to obtain ${{\hat{\delta}}_{{\rm{wf}}}}$, ${{\hat{\delta}}_{{\rm{if}}}}$, and ${{\hat{\delta}}_{{\rm{rnd}}}}$, respectively. According to (I), it is obvious that $\hat{\delta}\mathop{=}\limits^{Q\to+\infty}{\cal O}\left({{\sigma^{2}}{M^{2}}{Q^{-1}}}\right)$ holds for the three estimators. This completes the proof.